# The Role of Congeniality in Multiple Imputation for Doubly Robust Causal Estimation

### v1

@@ Abstract
This paper provides clear and practical guidance on the specification of imputation models when multiple imputation is used in conjunction with doubly robust estimation methods for causal inference. Through theoretical arguments and targeted simulations, we show that when a confounder has missing data the corresponding imputation model must include all variables used in either the propensity score model or the outcome model, and that these variables must appear in the same functional form as in the final analysis. Violating these conditions can lead to biased treatment effect estimates, even when both components of the doubly robust estimator are correctly specified. We present a mathematical framework for doubly robust estimation combined with multiple imputation, establish the theoretical requirements for proper imputation in this setting, and demonstrate the consequences of misspecification through simulation. Based on these findings, we offer concrete recommendations to ensure valid inference when using multiple imputation with doubly robust methods in applied causal analyses.
@@ 4.1 Primary Scenarios
Linear case with treatment heterogeneity: $Y=0.5X+Z_{C}+2Z_{P}+0.5X\times Z_{C}+\varepsilon$
Linear case without treatment heterogeneity: $Y=X+Z_{C}+2Z_{P}+\varepsilon$
Nonlinear case with treatment heterogeneity: $Y=0.5X+Z_{C}+2Z_{P}^{2}+0.5X\times Z_{C}+\varepsilon$
@@
$\mathbb{P}(M^{Y}=1)$ $=\left(1+\exp\{0.65+Z_{C}\}\right)^{-1}$
$\mathbb{P}(M^{Z_{C}}=1)$ $=\left(1+\exp\{1.15+0.5X\}\right)^{-1}$
@@ 4.2 Multiple Confounder Scenarios
For each scenario, we first generated two normally distributed variables: $Z_{C1}\sim N(1,1)$ and $Z_{C2}\sim N(1,1)$. We examined three data generation scenarios that varied the nonlinear relationships between $Z_{C2}$ and the exposure and/or outcome:
@@
$\mathbb{P}(M^{Y}=1)$ $=\left(1+\exp\{0.65+Z_{C1}\}\right)^{-1}$ (11)
$\mathbb{P}(M^{Z_{C1}}=1)$ $=\left(1+\exp\{1.15+0.5X\}\right)^{-1}$ (12)
@@ 4.3 Results
We conducted 500 simulation replications for each condition and computed the mean and standard deviation of the estimated treatment effects. Table LABEL:tbl-1 and Figure 4 show the correctly specified models as well as the impact of omitting $X$, $Y$, and $Z_{P}$ from the imputation models. Table LABEL:tbl-2 shows the impact of the true data generating mechanism having homogeneous treatment effects (instead of heterogeneous, as was the default). Table LABEL:tbl-3 shows the impact of incorrectly specifying nonlinear terms that were included in the analysis model in the imputation model.
Figure 4: Coverage probability by sample size for doubly robust estimates under different imputation specifications. Dashed line indicates nominal 95% coverage.
@@
Examining scenarios with multiple confounders where one was missing and the other had various nonlinear relationships with the exposure and outcome (Table LABEL:tbl-4LABEL:tbl-6), we find that imputation model sensitivity depends on whether the nonlinear relationship involves the exposure or the outcome. When the nonlinear relationship is between the additional confounder ($Z_{C2}$) and the exposure, the imputation model is not sensitive to this misspecification (assuming seperate imputation models are correctly specified by exposure status) (Table LABEL:tbl-4). However, when the nonlinear relationship is between the additional confounder and the outcome, incorrectly specifying $Z_{C2}$ in the imputation model as linear results in substantial bias (Table LABEL:tbl-5 and Table LABEL:tbl-6). Across all scenarios, the oversaturated imputation model recovers unbiased estimates provided the sample size is sufficient (i.e., it performs poorly when n = 100).
@@ 5.1 Data generation
# covariates
zp <- rnorm(n) # precision variable Z_P
zi <- rnorm(n) # auxiliary / additional covariate Z_I
zc <- rnorm(n, 1) # confounder Z_C
# treatment and nonlinear outcome with heterogeneity
@@ 5.2 Multiple imputation (separate by exposure)
# imputation formulas: include outcome and spline for c3
@@ 6 Conclusion
    Include all variables that appear in either the propensity score model or the outcome model in the imputation model, regardless of their causal role (confounder and precision variables).

### v2

@@ Abstract
This paper provides clear and practical guidance on the specification of imputation models when multiple imputation is used in conjunction with doubly robust estimation methods for causal inference. Through theoretical arguments and targeted simulations, we demonstrate that if a confounder has missing data, the corresponding imputation model must include all variables appearing in either the propensity score model or the outcome model, in addition to both the exposure and the outcome, and that these variables must enter the imputation model in the same functional form as in the final analysis. Violating these conditions can lead to biased treatment effect estimates, even when both components of the doubly robust estimator are correctly specified. We present a mathematical framework for doubly robust estimation combined with multiple imputation, establish the theoretical requirements for proper imputation in this setting, and demonstrate the consequences of misspecification through simulation. Based on these findings, we offer concrete recommendations to ensure valid inference when using multiple imputation with doubly robust methods in applied causal analyses.
@@ 4.1 Primary Scenarios
Linear case with treatment heterogeneity: $Y=0.5X+Z_{C}+2Z_{P}+0.5X\times Z_{C}+\varepsilon,$
Linear case without treatment heterogeneity: $Y=X+Z_{C}+2Z_{P}+\varepsilon,$
Nonlinear case with treatment heterogeneity: $Y=0.5X+Z_{C}+2Z_{P}^{2}+0.5X\times Z_{C}+\varepsilon.$
@@
$\mathbb{P}(M^{Y}=1)$ $=\left(1+\exp\{0.65+Z_{C}\}\right)^{-1},$
$\mathbb{P}(M^{Z_{C}}=1)$ $=\left(1+\exp\{1.15+0.5X\}\right)^{-1}.$
@@ 4.2 Multiple Confounder Scenarios
For each scenario, we first generated two normally distributed variables: $Z_{C1}\sim N(1,1)$ and $Z_{C2}\sim N(1,1)$. We examined three data generation scenarios that varied the nonlinear relationships between $Z_{C2}$ and the exposure and/or outcome.
@@
$\mathbb{P}(M^{Y}=1)$ $=\left(1+\exp\{0.65+Z_{C1}\}\right)^{-1},$
$\mathbb{P}(M^{Z_{C1}}=1)$ $=\left(1+\exp\{1.15+0.5X\}\right)^{-1},$
@@ 4.3 Results
We conducted 500 simulation replications for each condition and computed the mean and standard deviation of the estimated treatment effects. Table LABEL:tbl-1 and Figure 4 show the correctly specified models as well as the impact of omitting $X$, $Y$, $Z_{C}$, and $Z_{P}$ from the imputation models. Table LABEL:tbl-2 shows the impact of the true data generating mechanism having homogeneous treatment effects (instead of heterogeneous, as was the default). Table LABEL:tbl-3 shows the impact of incorrectly specifying nonlinear terms that were included in the analysis model in the imputation model.
Figure 4: Coverage probability by sample size for doubly robust estimates under different imputation specifications. Data were simulated under the scenario described as “Linear Data Generation (Treatment Heterogenity)” in Figure 2. The dashed line indicates nominal 95% coverage.
@@
Examining scenarios with multiple confounders where one was missing and the other had various nonlinear relationships with the exposure and outcome (Table LABEL:tbl-4, Table LABEL:tbl-5, Table LABEL:tbl-6), we find that imputation model sensitivity depends on whether the nonlinear relationship involves the exposure or the outcome. When the nonlinear relationship is between the additional confounder ($Z_{C2}$) and the exposure, the imputation model is not sensitive to this misspecification (assuming separate imputation models are correctly specified by exposure status) (Table LABEL:tbl-4). However, when the nonlinear relationship is between the additional confounder and the outcome, incorrectly specifying $Z_{C2}$ in the imputation model as linear results in substantial bias (Table LABEL:tbl-5 and Table LABEL:tbl-6). Across all scenarios, the oversaturated imputation model recovers unbiased estimates provided the sample size is sufficient (i.e., it performs poorly when n = 100).
@@ 5.1 Data generation
zp <- rnorm(n)
zi <- rnorm(n)
zc <- rnorm(n, 1)
@@ 5.2 Multiple imputation (separate by exposure)
# update imputation formulas to include spline
@@ 6 Conclusion
    Include all variables from the propensity score or outcome models in the imputation model, regardless of their causal role (confounders or precision variables), along with the exposure (when imputing a confounder or the outcome) and the outcome (when imputing a confounder).